2. Suppose that in your analysis of the market for cars, you discover that the Marginal Social Cost (MSC) of car production to differs from the Marginal Private Cost (MPC). While MPC = 6P-15, MSC = 6P - 33. Assume that MPB = MSB = 75-3P. a. Find the social equilibrium price and quantity for this market. b. Given the social equilibrium, is the private market producing too many or too few cars?
Answer #1

Ans) Market equilibrium ÷

MPB = MPC

75-3P = 6P-15

75+15 = 6P+3P

90 = 9P

P= 10

To find market quantity, plug the value of P in MPB equation

MPB = 75-3P = 75 - 3(10) = 75-30 = 45 units.

Social equilibrium ÷

MSB = MSC

75-3P = 6P-33

75+33 = 6P+3P

108 = 9P

P = 12

To find the social quantity, plug the value of P in MSB

75-3P = 75-3(12) = 75-36 = 39 units.

B) Since market quantity is more than social optimal quantity, market is over producing cars (too many).

That means, there exists negative externality, where marginal social cost is more than marginal private cost. And hence goods are over produced. To internalise this externality, government imposes tax equal to external cost. (External cost = MSC - MPC).